OpenStudy (anonymous):

use the quadratic formula to find the zeros of f. f(x) = x^2- 8x (smaller value & larger value)

OpenStudy (anonymous):

You know what the quadratic formula is? It's important!

OpenStudy (anonymous):

f(x)= x^2 -8x

OpenStudy (anonymous):

0=x^2-8x x(x-8)=0 x=0 and x=8